Will someone please show me how to find determinant A by expanding by column 1,
Given Matrix A:
(1 3 4)
(2 -1 3)
(3 1 -2)
Any help would be great.
Thanks, -Nick.

Answer by Alan3354(69443)

Given Matrix A:
(1 3 4)
(2 -1 3)
(3 1 -2)
--------
Using row 1:
Det = 1(2-3) - 3(-4-9) + 4(2+3)
= -1 + 39 + 20 = 58
-----------------
Using column 1:
Det = 1(2-3) - 2(-6-4) + 3(9+4)
= -1 + 20 + 39 = 58